A monopolist can produce with costs of MC=AC=30. The inverse demand curve for the product is p=90-q. A. If the firm adopts a uniform pricing strategy, what would be the price, output, consumer surplus, producer surplus, and deadweight loss? B. If the firm is able to perfectly price discriminate?
1
Expert's answer
2019-07-24T16:05:12-0400

If the firm adopts a uniform pricing strategy, then in equilibrium MR = MC:

MR = TR' = 90 - 2q,

90 - 2q = 30,

q = 30,

p = 90 - 30 = 60.

Consumer surplus CS = 0.5*(90 - 60)*30 = 450.

Producer surplus PS = 30*(60 - 30) = 900.

Deadweight loss DWL = 0.5*(60 - 30)*(60 - 30) = 450.